OpenStudy (anonymous):

when a polynomials p(x) is divided by x+3, the remainder is 2. which point must be on the graph of the corresponding function y=p(x)

ganeshie8 (ganeshie8):

Since \(p(x)\) leaves remainder \(2\) when divided by \(x + 3\) , so, \( p(x) = (x+3)\times p(x) + 2\)

ganeshie8 (ganeshie8):

plugin \(x = -3\), and see what you get

OpenStudy (anonymous):

oh I see! Thanks!